OpenStudy (anonymous):

VERIFY THE IDENTITY TAN^2X/SECX+1=1-COSX/COSX

OpenStudy (anonymous):

tan^2x = sec^2x-1 so now you have sec^2x-1 / secx+1 = 1-cosx/cosx

OpenStudy (anonymous):

from sin^2x + cos^2x = 1 you divide everything by cos^2x

OpenStudy (anonymous):

which gives you tan^2x + 1 = sec^2x